# Match-and-Fuse: Consistent Generation from Unstructured Image Sets

### v1

@@ A.2 Baselines
For IC-LoRA [lhhuang2024iclora], we use the most general Film Storyboard checkpoint with a LoRA weight of 0.8. IC-LoRA and FLUX [flux2024] are run with identical control signals as the method. FLUX and FLUX Kontext [labs2025flux1kontextflowmatching] share the same random seed across images, and FLUX Kontext is prompted by the instruction ”Turn the <<object>> into …”. For Edicho [bai2024edicho], we use the first image as a reference for the rest. Metrics in Tabs. 1 and 2 are computed with its public inpainting implementation available at publication, using the shared region masks from evaluation. The results in Fig. 9 with the ControlNet pipeline were provided by the authors.
@@ A.3 Evaluation Setup
CLIP score is computed with LongCLIP-L [zer0int_LongCLIP_L_Diffusers] and prompts $\mathcal{P}^{shared}$]. $\mathcal{P}_{i}^{non-shared}$].”.
The User Study evaluating consistency was conducted on 80 random edits from our benchmark using the Prolific [prolific2024] platform. Participants were shown the source image set, $\mathcal{P}^{shared}$, outputs from Match-and-Fuse and a baseline in a randomized 2AFC setup (see the interface in Fig. A.4). They answered: ”The source image set contains shared foreground object(s). In which image set, A or B, are these objects edited more consistently across images?”. We collected 1,600 judgments from 40 participants, with 5 respondents per comparison.
The automatic VLM evaluation was conducted on the full benchmark using GPT-5 [openai2025gpt5] with the system prompt: ”You will receive an image to evaluate along with an evaluation question. Output A or B.”. Each comparison used the same inputs as human participants: the compiled image (source set with randomized generations A and B), $\mathcal{P}^{shared}$, and the user-study question.
Our DINO-MatchSim metric quantifies fine-grained cross-image consistency in appearance and structure. Generated images are compared in the DINOv3 [simeoni2025dinov3] feature space at fixed correspondences $\textit{NN}_{ij}$—nearest neighbors from background-removed source images. DINOv3 is used for its strong and versatile representation and to ensure independence from the 2D matches used by our method. In practice, pairwise cosine similarities $S_{ij}$ (see Eq. 6) are averaged into $\bar{S}$, then transformed as: $\textit{DINO-MatchSim}=\exp((\bar{S}-1)/\tau)$ with $\tau=0.6$. This mapping amplifies consistency differences by compressing negative similarities into small scores ($\approx 0–0.2$) and expanding positive similarities into a wider interval ($\approx 0.2-1$).
@@
Subject-driven 131 357 CustomConcepts101 [kumari2022customdiffusion] DreamBooth [ruiz2023dreambooth]
Subject & Background 11 27 DyNeRF [li2022neural3dvideosynthesis] Instruct-NeRF2NeRF [instructnerf2023] Mip-NeRF [barron2021mipnerf] royalty-free videos
@@
Figure A.4: User study interface. Example of an interface with a question comparing Match-and-Fuse to FLUX [flux2024].
@@ Appendix B Runtime & Memory
During the extraction of 2D matches in preprocessing, RoMA has a runtime of 198.8 ms per image pair on an RTX6000 GPU [edstedt2023romarobustdensefeature]. In Fig. B.5, we report the per-image runtime of our full method as a function of $N$, progressively adding its components. As also discussed in Sec. 4.4, Match-and-Fuse maintains linear runtime for $N\geq 5$ by constraining graph connectivity to a node degree of 4. Since Feature Guidance noticeably increases runtime, it can be omitted when efficiency is critical, trading fine-grained consistency for speed.
@@ Appendix C Adapting Extended Attention for Our Task
We explore the additional qualitative baseline of prompting consistency through extended attention, widely used in UNets for cross-sample appearance sharing [alaluf2023crossimage, Cao_2023_ICCV, tokenflow2023, tewel2024training], and observe that in FLUX, this approach leads to artifacts and lacks semantic reasoning. We share our findings to inform future research.
@@
The positional embeddings (PEs) with a 2D token coordinate must be defined for all keys. [tewel2025addit] showed that in FLUX they strongly influence the spatial object placement. However, since extra keys are not truly present in the generated image plane, the choice of their positions is ambiguous. We test three different strategies: (i) Re-using $PE_{self}$. This variant yields identical generations that completely disregard the control signal (Fig. C.7a); (ii) Continuing positions analogous to PEs of spatially concatenated images in the grid generation. Despite this similarity, without a prompt indicating a joint canvas, this setting leads to artifacts and fails to achieve consistent outputs (Fig. C.7b); (iii) We extend only with keys that have a source 2D match and warp PEs of $K_{I^{self}}$ such that matched keys share the same position. Additionally, extra keys are scaled by $k$ to control their temperature in softmax (Fig. C.7c). With $k=1$, generations approach consistency at the cost of severe quality degradation. Reducing $k$ alleviates the artifacts, but compromises consistency before image quality is fully restored. In contrast, our method achieves consistency while avoiding artifacts (Fig. C.7d). These findings demonstrate a strong connection between features that share the same position in RoPE. Exploring this behavior may open future research directions.

### v2

@@ A.2 Baselines
For IC-LoRA [22], we use the most general Film Storyboard checkpoint with a LoRA weight of 0.8. IC-LoRA and FLUX [6] are run with identical control signals as the method. FLUX and FLUX Kontext [25] share the same random seed across images, and FLUX Kontext is prompted by the instruction ”Turn the <<object>> into …”. For Edicho [3], we use the first image as a reference for the rest. Metrics in Tabs. 1 and 2 are computed with its public inpainting implementation available at publication, using the shared region masks from evaluation. The results in Fig. 9 with the ControlNet pipeline were provided by the authors.
@@ A.3 Evaluation Setup
CLIP score is computed with LongCLIP-L [53] and prompts $\mathcal{P}^{shared}$]. $\mathcal{P}_{i}^{non-shared}$].”.
The User Study evaluating consistency was conducted on 80 random edits from our benchmark using the Prolific [32] platform. Participants were shown the source image set, $\mathcal{P}^{shared}$, outputs from Match-and-Fuse and a baseline in a randomized 2AFC setup (see the interface in Fig. A.4). They answered: ”The source image set contains shared foreground object(s). In which image set, A or B, are these objects edited more consistently across images?”. We collected 1,600 judgments from 40 participants, with 5 respondents per comparison.
The automatic VLM evaluation was conducted on the full benchmark using GPT-5 [31] with the system prompt: ”You will receive an image to evaluate along with an evaluation question. Output A or B.”. Each comparison used the same inputs as human participants: the compiled image (source set with randomized generations A and B), $\mathcal{P}^{shared}$, and the user-study question.
Our DINO-MatchSim metric quantifies fine-grained cross-image consistency in appearance and structure. Generated images are compared in the DINOv3 [39] feature space at fixed correspondences $\textit{NN}_{ij}$—nearest neighbors from background-removed source images. DINOv3 is used for its strong and versatile representation and to ensure independence from the 2D matches used by our method. In practice, pairwise cosine similarities $S_{ij}$ (see Eq. 6) are averaged into $\bar{S}$, then transformed as: $\textit{DINO-MatchSim}=\exp((\bar{S}-1)/\tau)$ with $\tau=0.6$. This mapping amplifies consistency differences by compressing negative similarities into small scores ($\approx 0–0.2$) and expanding positive similarities into a wider interval ($\approx 0.2-1$).
@@
Subject-driven 131 357 CustomConcepts101 [24] DreamBooth [34]
Subject & Background 11 27 DyNeRF [26] Instruct-NeRF2NeRF [18] Mip-NeRF [5] royalty-free videos
@@
Figure A.4: User study interface. Example of an interface with a question comparing Match-and-Fuse to FLUX [6].
@@ Appendix B Runtime & Memory
During the extraction of 2D matches in preprocessing, RoMA has a runtime of 198.8 ms per image pair on an RTX6000 GPU [12]. In Fig. B.5, we report the per-image runtime of our full method as a function of $N$, progressively adding its components. As also discussed in Sec. 4.4, Match-and-Fuse maintains linear runtime for $N\geq 5$ by constraining graph connectivity to a node degree of 4. Since Feature Guidance noticeably increases runtime, it can be omitted when efficiency is critical, trading fine-grained consistency for speed.
@@ Appendix C Adapting Extended Attention for Our Task
We explore the additional qualitative baseline of prompting consistency through extended attention, widely used in UNets for cross-sample appearance sharing [1, 8, 17, 42], and observe that in FLUX, this approach leads to artifacts and lacks semantic reasoning. We share our findings to inform future research.
@@
The positional embeddings (PEs) with a 2D token coordinate must be defined for all keys. [41] showed that in FLUX they strongly influence the spatial object placement. However, since extra keys are not truly present in the generated image plane, the choice of their positions is ambiguous. We test three different strategies: (i) Re-using $PE_{self}$. This variant yields identical generations that completely disregard the control signal (Fig. C.7a); (ii) Continuing positions analogous to PEs of spatially concatenated images in the grid generation. Despite this similarity, without a prompt indicating a joint canvas, this setting leads to artifacts and fails to achieve consistent outputs (Fig. C.7b); (iii) We extend only with keys that have a source 2D match and warp PEs of $K_{I^{self}}$ such that matched keys share the same position. Additionally, extra keys are scaled by $k$ to control their temperature in softmax (Fig. C.7c). With $k=1$, generations approach consistency at the cost of severe quality degradation. Reducing $k$ alleviates the artifacts, but compromises consistency before image quality is fully restored. In contrast, our method achieves consistency while avoiding artifacts (Fig. C.7d). These findings demonstrate a strong connection between features that share the same position in RoPE. Exploring this behavior may open future research directions.
Appendix D Downstream Applications
In addition to the extended applications in Sec. 3.3, in Fig. D.8 we demonstrate a downstream video editing application enabled by our method and not achievable with image-level editing alone, further motivating the task of consistent set-to-set editing.
Figure D.8: Additional experiments: multi-cut ad editing. Product advertisements often depict multiple cuts of the same object in different contexts rather than one continuous shot, which can be obtained by conditioning an I2V model [44] on keyframes. Our method enables infinite coherent edits with cross-cut consistency, where one-to-one methods fail.